OpenStudy (anonymous):

find the derivative of (6 + x)/(1 - 6x)

OpenStudy (turingtest):

quotient rule:\[\frac d{dx}(\frac{f(x)}{g(x)})=\frac{f'(x)\cdot g(x)-g'(x)\cdot f(x)}{[g(x)]^2}\]

OpenStudy (anonymous):

i dont know how to use the rule that well..could you help

OpenStudy (turingtest):

call\[f(x)=6+x\]and\[g(x)=1-6x\]find\[f'(x)\]and\[g'(x)\]

OpenStudy (anonymous):

37/(1-6x)^2